# MoMuSE: Momentum Multi-modal Target Speaker Extraction for Real-time Scenarios with Impaired Visual Cues

### v1

@@ Abstract
Audio-visual Target Speaker Extraction (AV-TSE) aims to isolate the speech of a specific target speaker from an audio mixture using time-synchronized visual cues. In real-world scenarios, visual cues are not always available due to various impairments, which undermines the stability of AV-TSE. Despite this challenge, humans can maintain attentional momentum over time, even when the target speaker is not visible. In this paper, we introduce the Momentum Multi-modal target Speaker Extraction (MoMuSE), which retains a speaker identity momentum in memory, enabling the model to continuously track the target speaker. Designed for real-time inference, MoMuSE extracts the current speech window with guidance from both visual cues and dynamically updated speaker momentum. Experimental results demonstrate that MoMuSE exhibits significant improvement, particularly in scenarios with severe impairment of visual cues.  [^1] Demo: https://mrjunjieli.github.io/demo_page/MoMuSE/index.html.
Code: https://github.com/mrjunjieli/MoMuSE.
@@
Index Terms—  Audio-visual, Momentum, Multi-modal, Target Speaker Extraction, Visual Impairments
1 Introduction
@@
Audio-visual Target Speaker Extraction (AV-TSE) [1, 2, 3, 4] aims to extract the speech of target speaker with visual cues, such as lip or face image sequences, as references. Visual cues, known for their resilience against acoustic noise, have demonstrated superior performance compared to audio-only approaches [5, 6, 7] in complex acoustic environments.
@@
Despite the significant benefits that visual information provides to the AV-TSE task, challenges arise when visual cues are temporarily unavailable in real-world scenarios, such as pose variations or the absence of the target person’s video [8]. Additionally, issues like lip occlusion or a low-quality camera [9] can result in unclear lip areas. These challenges dynamically disrupt the reliance of AV-TSE on time-aligned cue sequences.
@@
Several related studies have tried to address the aforementioned challenges. Sadeghi and Alameda-Pineda [10, 11] proposed switching from an audio-visual variational auto-encoder (VAE) to an audio-only VAE when visual quality is poor. Wu et al. [12] incorporated an attention mechanism, selecting relevant visual features based on mixed audio features. In contrast to discarding low-quality visual features, Pan et al. [13] adopted an innovative approach by reconstructing corrupted video through audio-visual correspondence. Furthermore, the VS model [9] and the audio-visual SpeakerBeam [14, 15] learn a speaker embedding from pre-enrolled speech and rely on this embedding when visual cues are unreliable. Liu et al. [16] propose a triple loss function, where visual cues are leveraged only during the training phase and are not directly fused into the model’s input features.
@@
However, the aforementioned approaches introduce an additional step of using pre-enrollment, and the pre-enrolled speech may not align with the actual acoustic scenario during usage. Instead of pre-enrollment, we propose tracking speaker identity using speech derived from previous time steps, offering a more accurate representation of the speaker’s current vocal characteristics. While leveraging such “self-enrollment” speech for second-phase inference has proven effective in tasks like speech enhancement [17], audio-based speech separation [18, 19, 20, 21], and target speaker extraction [22, 23, 24, 25], the performance degrades if the self-enrollment speech is of poor quality.
@@
In this paper, we propose the Momentum Multi-modal target Speaker Extraction (MoMuSE), a novel framework that employs a memory bank with a momentum mechanism to dynamically track and update the speaker’s identity using previous time steps, This dynamic updating ensures that the memory bank retains the most accurate speaker embedding, enabling MoMuSE to maintain focus on the target speaker during real-time inference, even when visual cues are corrupted or unavailable. Experiments demonstrate that our proposed MoMuSE achieves a momentum effect, enabling target speaker extraction to continue functioning even when visual cues are completely absent.
2 Visual Impairments
Fig. 1: Examples of normal and impaired visual frames.
@@ 2 Visual Impairments
Fig. 2: (a) Overall structure of MoMuSE. Modules in gray denote the original structures from MuSE, while modules in other colors denote the proposed new structures. (b) Speaker Extractor. (c) The detailed structure of ASEU, which is an attention based module. $\otimes$, $\ominus$ and $\oplus$ refer to point-wise multiplication, concatenation and point-wise addition.
@@
3 Proposed Method
To enhance the stability of AV-TSE systems in the aforementioned impaired scenarios and avoid introducing a pre-enrollment process [26, 27, 28, 29], we aim to leverage the advantages of dynamic visual frames when visual cues are reliable while incorporating alternative speaker voiceprint cue when visual information is impaired. In this paper, we introduce MoMuSE, a novel framework that employs a dynamically updated speaker momentum as an additional reference. MoMuSE enables robust speaker extraction even in the presence of visual impairments, presenting a more effective way to harness the complementary nature of visual and audio modalities.
3.1 Recap of MuSE
@@
MuSE consists of four main modules: an audio encoder, an audio decoder, a visual encoder, and a speaker extractor. The audio encoder transforms the time-domain mixed speech $\mathbf{y}$ into a latent representation $\mathbf{Y}$, and the audio decoder reconstructs the extracted target speech $\mathbf{x}$ from its latent representation $\hat{\mathbf{X}}$. The visual encoder processes the visual sequence $\mathbf{v}$ to generate visual embeddings $\mathbf{V}$. The speaker extractor iteratively refines the quality of the extracted speech using R extractor blocks. Each extractor block includes a speaker encoder and a mask estimator, as illustrated in Fig 2 (b). The speaker encoder fuses the estimated speech representation $\hat{\mathbf{X}}^{r-1}$ and the visual representation $\mathbf{V}$ to generate a speaker embedding $\mathbf{e}^{r}_{c}$. The mask estimator then generates a mask $\mathbf{M}^{r}$ to isolate the target speech. Here, $r=(1,2,\ldots,R)$ indicates the index of the extractor block, with distinct parameter weights for each block.
@@
3.2 MoMuSE architecture
MoMuSE shares a similar structure with MuSE, yet it incorporates a memory bank that allows it to leverage historical voiceprint information from previously extracted speech, as illustrated in Fig 2 (a). The memory bank comprises R anchor speaker embeddings, denoted as ${\mathbf{e}^{1}_{a},\dots,\mathbf{e}^{R}_{a}}$, which store the historical voiceprint information of the target speaker. Here, the subscript $a$ denotes “anchor”. In addition, to effectively combine historical and current speaker information, we propose an Anchor Speaker Embedding Updating (ASEU) sub-module within each speaker extractor block. This sub-module integrates historical embeddings into the extraction process for the current window, adapting based on the visual features’ quality in the current window. We implement the ASEU sub-module using an additive attention fusion mechanism inspired by [14, 31] [^2] https://github.com/sooftware/attentions/blob/master/attentions.py, as shown in Fig 2 (c).
To maintain temporal resolution consistency, we apply temporal replication to the speaker embeddings [^3] For notational clarity, the superscript $r$ denoting the $r$-th block will be neglected in Section 3.2 and Section 3.3.
$\mathbf{E}_{\psi}\in\mathbb{R}^{H\times L}=\text{Repeat}(\mathbf{e}_{\psi}\in\mathbb{R}^{H\times 1})$ (1)
@@ 3.2 MoMuSE architecture
$\mathbf{s}_{\psi}$ $=\text{Linear}(\text{Tanh}(\text{Linear}(\mathbf{E}_{\psi})+\text{Linear}(\mathbf{V})))$ (2)
where $\mathbf{V}\in\mathbb{R}^{H\times L}$ represents the visual feature. These score vectors are then normalized using a softmax function to produce attention weights:
$\mathbf{a}_{\psi,l}=\frac{\exp(\mathbf{s}_{\psi,l})}{\exp(\mathbf{s}_{c,l})+\exp(\mathbf{s}_{a,l})}\quad(l=1,2,\ldots,L),$ (3)
@@
$\mathbf{E}_{m}=\mathbf{a}_{c}\otimes\mathbf{E}_{c}+\mathbf{a}_{a}\otimes\mathbf{E}_{a},$ (4)
@@
3.3 Momentum mechanism
In online real-time processing [23], the model begins processing only after the incoming mixed input speech accumulates to a predefined initialization length $L_{\text{init}}$ (in seconds). Once this initialization is complete, the model processes the mixed speech $y$ using a window size of $L_{\text{win}}$ (in seconds), and then advances forward by a step size of $L_{\text{shift}}$ (in seconds). In subsequent discussions, we refer to the $t$-th processing window as the $(t)$-th window step. At the $t$-th window step, the model takes the mixed speech $\mathbf{y}(t)$ and the corresponding visual sequence $\mathbf{v}(t)$ as inputs, and outputs the estimated target speech $\hat{\mathbf{x}}(t)$:
$\hat{\mathbf{x}}(t)=\text{MoMuSE}(\mathbf{y}(t),\mathbf{v}(t))$ (5)
@@ 3.3 Momentum mechanism
Initialization: To maintain continuous attention on the target speaker in unreliable visual scenarios, the historical voiceprint information of the target speaker needs to be preserved. We introduce a memory bank to store this information. At the first window step (initialization), the memory bank is empty. The dimension of the mixed speech of current window is $\mathbf{y}(1)\in\mathbb{R}^{1\times L_{\text{init}}}$. Given that the memory bank is initially empty, the ASEU sub-module is not utilized at this step. Instead, the current speaker embedding sequence $\mathbf{E}_{c}(1)$ is directly used as input to the mask estimator. Once the corresponding target speech is estimated, the memory bank is initialized with the current speaker embedding $\mathbf{e}_{c}(1)$:
@@
where $\overline{\mathbf{a}}_{c}(t)\in\mathbb{R}^{1}$ denotes the mean of the attention weights $\mathbf{a}_{c}(t)$ across the time dimension, and $\theta$ is a predefined threshold.
3.4 Loss function
@@
$\mathcal{L}_{\text{SI-SNR}}(\mathbf{x},\hat{\mathbf{x}})=-10\log_{10}\frac{||\frac{<\hat{\mathbf{x}},\mathbf{x}>\mathbf{x}}{||\mathbf{x}||^{2}}||^{2}}{||\hat{\mathbf{x}}-\frac{<\hat{\mathbf{x}},\mathbf{x}>\mathbf{x}}{||\mathbf{x}||^{2}}||},$ (8)
$\mathcal{L}_{\text{CE}}(\mathbf{e}^{r}_{c})=-\sum^{N}_{n=1}d_{n}\log(\operatorname{softmax}(\mathbf{W}^{r}\mathbf{e}^{r}_{c})),$ (9)
where $d_{n}$ is class label for speaker $n$ and $\mathbf{x}$ is target speech. $N$ is the number of speakers in dataset. $\mathbf{W}^{r}$ is a learnable weight matrix in $r$-th block used to project speaker embedding $\mathbf{e}^{r}_{c}$ to a one-hot class label.
@@
$\mathcal{L}_{\text{pe}}=\sum^{R}_{r=1}\frac{||\mathbf{a}^{r}_{a}||_{1}}{L}.$ (10)
@@
3.5 Training procedure
In autoregressive training, the past extracted speech is used as input for the current extraction, which can significantly increase computational overhead if every sliding window is optimized. Drawing inspiration from the training settings in NeuroHeed [23], we propose a Segment-Level Optimization (Seg) strategy that simulates only the first two window steps during the training phase. This approach effectively reduces computational cost while still capturing the essence of the autoregressive process.
@@ 3.5 Training procedure
Utt: MoMuSE processes the whole utterance of visual and audio inputs, ensuring the model encounters long data, which aims to improve the robustness of the speaker encoder:
@@
Seg: To simulate autoregressive training, the entire utterance is split into two segments. The extraction of the second segment incorporates past extracted speaker embedding as input. Following the settings in NeuroHeed [23], we define the window length $L_{\text{win}}\sim\mathcal{U}(1.05s,3.2s)$, shift length $L_{\text{shift}}\sim\mathcal{U}(0.05s,0.2s)$, and initialization length $L_{\text{init}}\sim\mathcal{U}(0.05s,3s)$, where $\mathcal{U}$ denotes a uniform distribution, and the unit is in seconds (s). In the first window step, MoMuSE processes the mixed audio and visual sequence from the first segment:
$\hat{\mathbf{x}}(1),\mathbf{e}^{r}_{c}(1)=\text{MoMuSE}(\mathbf{y}(1),\mathbf{v}(1)).$ (13)
@@
$\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{SI-SNR}}(\mathbf{x}(2),\hat{\mathbf{x}}(2))+\lambda\sum^{R}_{r=1}\mathcal{L}_{\text{CE}}(\overline{\mathbf{E}}^{r}_{m}(2)),$ (15)
@@
$\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{utt}}+\beta\mathcal{L}_{\text{seg}}+\gamma\mathcal{L}_{\text{pe}},$ (16)
@@
4 Experimental Details
4.1 Dataset

### v2

# Junjie Li1, Ke Zhang2, Shuai Wang2,3,∗, Kong Aik Lee1,∗, Man-Wai MAK1, Haizhou Li2,3

@@ Abstract
Audio-visual Target Speaker Extraction (AV-TSE) aims to isolate the speech of a specific target speaker from an audio mixture using time-synchronized visual cues. In real-world scenarios, visual cues are not always available due to various impairments, which undermines the stability of AV-TSE. Despite this challenge, humans can maintain attentional momentum over time, even when the target speaker is not visible. In this paper, we introduce the Momentum Multi-modal target Speaker Extraction (MoMuSE), which retains a speaker identity momentum in memory, enabling the model to continuously track the target speaker. Designed for real-time inference, MoMuSE extracts the current speech window with guidance from both visual cues and dynamically updated speaker momentum. Experimental results demonstrate that MoMuSE exhibits significant improvement, particularly in scenarios with severe impairment of visual cues.
Index Terms:
@@
Audio-visual, Momentum, Multi-modal, Target Speaker Extraction, Visual Impairments
I Introduction
@@
Audio-visual Target Speaker Extraction (AV-TSE) [1, 2, 3] aims to extract the speech of a target speaker using visual cues, such as lip or face image sequences, as references. AV-TSE, known for their resilience against acoustic noise, have demonstrated superior performance compared to audio-only approaches [4, 5] in complex acoustic environments.
@@
Despite the significant benefits that visual information provides to AV-TSE tasks, challenges arise when visual cues are temporarily unavailable in real scenarios, such as absence of the target person’s video. Additionally, issues like lip occlusion or a low-quality camera [6] can result in unclear lip areas. These challenges make AV-TSE systems relying on time-aligned sequences impractical under real-world environments.
@@
Several related studies have tried to address the aforementioned challenges. Sadeghi and Alameda-Pineda [7, 8] propose switching from an audio-visual variational auto-encoder (VAE) to an audio-only VAE when visual quality is poor. Wu et al. [9] incorporate an attention mechanism, selecting relevant visual features based on mixed audio features. In contrast to discarding low-quality visual features, Pan et al. [10] adopt an innovative approach by reconstructing corrupted video through audio-visual correspondence. Furthermore, the VS model [6] and the audio-visual SpeakerBeam [11, 12] learn a speaker embedding from pre-enrolled speech and rely on this embedding when visual cues are unreliable. Liu et al. [13] propose a triple loss function, where visual cues are leveraged only during the training phase and are not directly fused into the model’s input features.
@@
Despite achieving good performance, above approaches introduce an additional pre-enrollment step, and the acoustic environments of the pre-enrolled speech may not match the deployment environments. Instead of pre-enrollment, we propose tracking speaker identity using speech derived from previous time steps, offering a more accurate representation of the speaker’s current vocal characteristics. While leveraging such “self-enrollment” speech for second-phase inference has proven effective in tasks like speech enhancement [14], audio-based speech separation [15, 16, 17, 18], and target speaker extraction [19, 20, 21], the performance degrades if the self-enrollment speech is of poor quality.
@@
In this paper, we propose Momentum Multi-modal target Speaker Extraction (MoMuSE)  [^1] Demo: https://mrjunjieli.github.io/demo_page/MoMuSE/index.html, a novel framework that employs a memory bank with a momentum mechanism to dynamically track and update the speaker’s identity using previous time steps, This dynamic updating ensures that the memory bank retains the most accurate speaker embedding, enabling MoMuSE to maintain focus on the target speaker during real-time, even when visual cues are corrupted. Experiments demonstrate that our proposed MoMuSE achieves a momentum effect, enabling TSE to continue functioning even when visual cues are completely absent.
II Visual Impairments
Figure 1: Examples of normal and impaired visual frames.
@@ II Visual Impairments
Figure 2: (a) Overall structure of MoMuSE. Modules in gray denote the original structures from MuSE, while modules in other colors denote the proposed new structures. (b) Speaker Extractor. (c) The detailed structure of Anchor Speaker Embedding Updating (ASEU), which is an attention based module. $\otimes$, $\ominus$ and $\oplus$ refer to point-wise multiplication, concatenation and point-wise addition.
@@
III Proposed Method
To enhance the stability of AV-TSE systems in the visual impairment scenarios and avoid introducing a pre-enrollment process [22, 23, 24], we leverage the advantages of dynamic visual frames when visual cues are reliable while incorporating alternative speaker voiceprint cue when visual information is impaired. In this paper, we introduce MoMuSE, a novel framework that employs a dynamically updated speaker momentum as an additional reference. MoMuSE enables robust speaker extraction even in the presence of visual impairments, presenting a more effective way to harness the complementary nature of visual and audio modalities.
III-A Recap of MuSE
@@
MuSE comprises four key modules: an audio encoder, an audio decoder, a visual encoder, and a speaker extractor. The audio encoder converts the time-domain mixed speech $\mathbf{y}$ into a latent representation $\mathbf{Y}$, and the audio decoder reconstructs the extracted target speech $\mathbf{x}$ from its latent representation $\hat{\mathbf{X}}$. The visual encoder processes the visual sequence $\mathbf{v}$ to generate visual embeddings $\mathbf{V}$. The speaker extractor iteratively refines the quality of the extracted speech using $R$ extractor blocks. Each block includes a speaker encoder and a mask estimator, as illustrated in Fig 2 (b). The speaker encoder fuses the estimated speech representation $\hat{\mathbf{X}}^{r-1}$ and the visual representation $\mathbf{V}$ to generate a speaker embedding $\mathbf{e}^{r}_{c}$. The mask estimator then generates a mask $\mathbf{M}^{r}$ to isolate the target speech. Here, $r=(1,2,\ldots,R)$ indicates the index of the extractor block, with distinct parameter weights for each block.
@@
III-B MoMuSE Architecture
MoMuSE shares a similar structure with MuSE, yet it incorporates a memory bank that allows it to leverage historical voiceprint information from previously extracted speech, as illustrated in Fig 2(a). The memory bank comprises $R$ anchor speaker embeddings, denoted as ${\mathbf{e}^{1}_{a},\dots,\mathbf{e}^{R}_{a}}$, which store the historical voiceprint information of the target speaker. Here, the subscript $a$ denotes “anchor”. In addition, to effectively combine historical and current speaker information, we propose an Anchor Speaker Embedding Updating (ASEU) sub-module within each speaker extractor block. This sub-module integrates historical embeddings into the extraction process for the current window, adapting based on the visual features’ quality in the current window. We implement the ASEU sub-module using an additive attention fusion [^2] https://github.com/sooftware/attentions/blob/master/attentions.py mechanism inspired by [11, 26], as shown in Fig 2(c).
To maintain temporal resolution consistency, we apply temporal replication to the speaker embeddings [^3] For notational clarity, the superscript $r$ denoting the $r$-th block will be neglected in Section III-B and Section III-C.
$\mathbf{E}_{\psi}\in\mathbb{R}^{H\times L}=\text{Repeat}(\mathbf{e}_{\psi}\in\mathbb{R}^{H\times 1}),$ (1)
@@ III-B MoMuSE Architecture
$\mathbf{s}_{\psi}\in\mathbb{R}^{1\times L}$ $=\text{Linear}(\text{Tanh}(\text{Linear}(\mathbf{E}_{\psi})+\text{Linear}(\mathbf{V}))),$ (2)
where $\mathbf{V}\in\mathbb{R}^{H\times L}$ represents the visual feature. These score vectors are then normalized using a softmax function to produce element-wise attention weights:
$\mathbf{a}_{\psi,l}\in\mathbb{R}^{1}=\frac{\exp(\mathbf{s}_{\psi,l})}{\exp(\mathbf{s}_{c,l})+\exp(\mathbf{s}_{a,l})}\quad(l=1,2,\ldots,L),$ (3)
@@
$\mathbf{E}_{m}\in\mathbb{R}^{H\times L}=\mathbf{a}_{c}\otimes\mathbf{E}_{c}+\mathbf{a}_{a}\otimes\mathbf{E}_{a},$ (4)
@@
III-C Momentum Mechanism
[…]
$\hat{\mathbf{x}}(t)=\text{MoMuSE}(\mathbf{y}(t),\mathbf{v}(t)).$ (5)
@@ III-C Momentum Mechanism
Initialization: To maintain continuous attention on the target speaker in unreliable visual scenarios, the historical voiceprint information of the target speaker needs to be preserved. We introduce a memory bank to store this information. At the first window step (initialization), the memory bank is empty. The dimension of the mixed speech of the current window is $\mathbf{y}(1)\in\mathbb{R}^{1\times L_{\text{init}}}$. Given that the memory bank is initially empty, the ASEU sub-module is not utilized at this step. Instead, the current speaker embedding sequence $\mathbf{E}_{c}(1)$ is directly used as input to the mask estimator. Once the corresponding target speech is estimated, the memory bank is initialized with the current speaker embedding $\mathbf{e}_{c}(1)$:
@@
where $\overline{\mathbf{a}}_{c}(t)\in\mathbb{R}^{1}$ is the mean of the attention weights $\mathbf{a}_{c}(t)$ across the time dimension, and $\theta$ is a predefined threshold.
III-D Loss Function
@@
$\mathcal{L}_{\text{SI-SNR}}(\mathbf{x},\hat{\mathbf{x}})=-10\log_{10}\frac{||\frac{<\hat{\mathbf{x}},\mathbf{x}>\mathbf{x}}{||\mathbf{x}||^{2}}||^{2}}{||\hat{\mathbf{x}}-\frac{<\hat{\mathbf{x}},\mathbf{x}>\mathbf{x}}{||\mathbf{x}||^{2}}||},\vspace{-1mm}$ (8)
$\mathcal{L}_{\text{CE}}(\mathbf{e}^{r}_{c})=-\sum^{N}_{n=1}d_{n}\log(\operatorname{softmax}(\mathbf{W}^{r}\mathbf{e}^{r}_{c})),\vspace{-1mm}$ (9)
where $d_{n}$ is class label for speaker $n$ and $\mathbf{x}$ is target speech. $N$ is the number of training speakers. $\mathbf{W}^{r}$ is a learnable weight matrix in the $r$-th block used to project speaker embedding $\mathbf{e}^{r}_{c}$ to a one-hot class label.
@@
$\mathcal{L}_{\text{pe}}=\sum^{R}_{r=1}\frac{||\mathbf{a}^{r}_{a}||_{1}}{L}.\vspace{-1mm}$ (10)
@@
III-E Training Procedure
In autoregressive training, the past extracted speech is used as input for the current extraction, which can significantly increase computational overhead if every sliding window is optimized. Inspired by the training settings in NeuroHeed [20], we propose a Segment-Level Optimization (Seg) strategy that simulates only the first two window steps during training. This approach effectively reduces computational cost while still capturing the essence of the autoregressive process.
@@ III-E Training Procedure
Utt: MoMuSE processes the whole utterance of visual and audio inputs, ensuring the model encounters long utterance, which aims to improve the robustness of the speaker encoder:
@@
Seg: To simulate autoregressive training, the entire utterance is split into two segments. The extraction of the second segment incorporates past extracted speaker embedding as input. Following the settings in NeuroHeed [20], we define the window length $L_{\text{win}}\sim\mathcal{U}(1.05\text{s},3.2\text{s})$, shift length $L_{\text{shift}}\sim\mathcal{U}(0.05\text{s},0.2\text{s})$, and initialization length $L_{\text{init}}\sim\mathcal{U}(0.05\text{s},3\text{s})$, where $\mathcal{U}$ denotes a uniform distribution, and the unit is in seconds (s). In the first window step, MoMuSE processes the mixed audio and visual sequence from the first segment:
$\hat{\mathbf{x}}(1),\mathbf{e}^{r}_{c}(1)=\text{MoMuSE}(\mathbf{y}(1),\mathbf{v}(1)).\vspace{-1mm}$ (13)
@@
$\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{SI-SNR}}(\mathbf{x}(2),\hat{\mathbf{x}}(2))+\lambda\sum^{R}_{r=1}\mathcal{L}_{\text{CE}}(\overline{\mathbf{E}}^{r}_{m}(2)),\vspace{-1mm}$ (15)
@@
$\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{utt}}+\beta\mathcal{L}_{\text{seg}}+\gamma\mathcal{L}_{\text{pe}},\vspace{-1mm}$ (16)
@@
IV Experimental Details
IV-A Dataset